I need to find the point slope form of the line with the following coordinates (5, -3), (0, -11).
This is what i have so far
y-(-3)= 8/5(x-5)
y+3 = 8/5x -8
is that correct

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

but why did you stop there?
keep going.
y+3 = 8/5x -8
subtract the 3
and you are finished
y=8x/5-5
standard form would be
5y=8x-5
5=8x-5y
8x-5y=5

Answer by Alan3354(69443)

I need to find the point slope form of the line with the following coordinates (5,-3), (0,-11).
This is what i have so far
y-(-3)= 8/5(x-5)
y+3 = 8/5x -8
-------------
One more step:
y = (8/5)x - 11
It's the slope-intercept form, actually.